# No Other Representation Component Is Needed: Diffusion Transformers Can Provide Representation Guidance by Themselves

### v4

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can both accelerate generative training and enhance the generation quality of diffusion transformers. However, existing approaches necessitate to either introduce an external and complex representation training framework or rely on a large-scale, pre-trained representation model to provide representation guidance during the original generative training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We therefore propose Self-Representation Alignment (SRA), a simple yet effective method that obtains representation guidance through a self-distillation manner. Specifically, SRA aligns the latent representation of the diffusion transformer in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the overall representation learning during only the generative training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements. Moreover, SRA not only largely outperforms approaches relying on auxiliary, complex representation training frameworks but also achieves performance comparable to methods that are heavily dependent on an external pre-trained representation model. Refer to caption Figure 1: Left: Methods like MaskDiT [84] and SD-DiT [86] use an external representation task to guide diffusion transformer. Middle: Methods like REPA [81] leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
@@ 1 Introduction
Diffusion transformers [62, 56, 10] and vision transformers [20, 52, 75] have held the dominant positions in visual generation and representation because of their scalability during pre-training [5, 22, 61, 72] and generalization capacity for downstream tasks [41, 50, 38, 48].
@@
Recently, many works [84, 86, 81, 42] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the training paradigm in representation learning (e.g., MAE’s [29], IBOT’s [85]) shown in Figure 1a or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [61], CLIP [65]) shown in Figure 1b to give representation guidance for the diffusion transformer during the original generative training. However, the former method requires a complex training framework design, and the latter method relies on a powerful model, which is trained on a massive amount of data with hundreds of GPUs. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance only during the generative training without external representation components?
@@
Figure 2: We empirically investigate the representations in diffusion transformers across different blocks and timesteps with the original SiT-XL/2 checkpoint trained for 7M iterations. Left: Using PCA [1] to visualize the latent features in SiT, we observe that the features lead a process from coarse to fine when increasing block layers and decreasing noise level. Right: A similar trend can also be seen in the linear probing results on ImageNet. Investigation of DiT is provided in Appendix A, which leads to the similar results as SiT.
@@
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a coarse to fine process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [56, 62]. As shown in Figure 2a, we first find out that the latent features in the diffusion transformer are progressively refined, moving from coarse to fine, as block layers increase and noise level decreases. Next, akin to the results in previous studies [81, 77], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2b. Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a coarse-to-fine discriminative process when only generative training is performed.
This distinctive trend motivates us to align the weaker representations in the diffusion transformer to the better ones in the original generative training, thereby enhancing the representation learning of the model without involving any external representation component. However, this approach is not straightforward during training. As observed in previous studies [12, 28], directly using the output of the same model as a supervision signal may lead to shortcut learning [26], which can compromise performance or even cause training collapse.
@@
Our approach: To overcome the abovementioned hurdles, we present Self-Representation Alignment (SRA), a simple but effective technique built on recent diffusion transformer architectures [56, 62]. As shown in Figure 1c, SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned on higher noise to that in later layer conditioned on lower noise to achieve self-representation alignment. Meanwhile, in order to make the training process more stable, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architect modification.
@@
Finally, we conduct comprehensive experiments to evaluate the effect of SRA. After a series of component-wise analyses within the design space, we show that SRA brings significant performance improvements to both DiTs [62] and SiTs [56]. Moreover, our ablation study highlights the crucial role of internal representations in the success of SRA, which supports our central hypothesis: diffusion transformers can achieve representation alignment without external components.
@@
  • We analyze the latent representation in the diffusion transformers and assume that the unique coarse to fine discriminative process makes it possible to achieve representation alignment without external components.
@@ 2.1 Preliminary: Training Object of DiT and SiT
As our method is built upon the denoise-based model (DiT) and flow-based model (SiT), to facilitate a more seamless introduction of SRA in the subsequent part, we now present a brief overview of the training object of these two types of models. We omit the class-condition here for simplicity because it can be easily added to all analogs without causing any changes to the mathematical equations. We leave more detailed mathematical descriptions of these types of models in Appendix B.
@@
$\mathcal{L}_{\text{simple}}=\mathbb{E}{{\mathbf{x}}_{t},\bm{\epsilon},t}\Big{[}||\bm{\epsilon}-\bm{\epsilon}_{\bm{\zeta}}({\mathbf{x}}_{t},t)||_{2}^{2}\Big{]}.$ (1)
@@
Where the marginal distribution of this ODE at time $p_{t}({\mathbf{x}})$ matches the marginal distribution of the forward process. To learn the velocity field, the model ${\mathbf{v}}_{\zeta}({\mathbf{x}}_{t},t)$ is trained to minimize the following loss function:
$\mathcal{L}_{\text{velocity}}=\mathbb{E}_{{\mathbf{x}}_{0},\bm{\epsilon},t}\big{[}||{\mathbf{v}}_{\bm{\zeta}}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{0}-\dot{\sigma}_{t}\bm{{\mathbf{\epsilon}}}||^{2}\big{]}.$ (4)
@@ 2.2 Self-Representation Alignment
Previous studies have demonstrated that learning good internal representation can both speed up the diffusion transformer’s training and improve the quality of its generated samples. In SRA, our insight is aligning the student’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on the lower noise of the teacher to conduct self-representation alignment without requiring any external representation components. As depicted in Figure 3, the goal of our simple training framework is to let the diffusion transformer not only predict noise-invariant but also align with better visual representations from itself. This operation thereby provides a simple way to enhance the representation learning in the diffusion transformer during generative training without the need to design complex representation regulation or introduce external representation foundation models.
Formally, let $f$ be the trainable student model and $f_{\ast}$ be the teacher model. Considering input noise latent, timestep, and condition to be ${\mathbf{x}}_{t}$, $t$, and $c$. Then, we can obtain the student encoder latent output ${\mathbf{y}}=f^{m}({\mathbf{x}}_{t},t,c)\in\mathbb{R}^{B\times N\times D}$, where $B,N,D>0$ are the batchsize, number of patches and the embedding dimension for $f$, and $m$ denote the output from the $m^{th}$ layer in $f$. Similarly, the output of the teacher can be expressed as ${\mathbf{y}}_{\ast}=f_{\ast}^{n}({\mathbf{x}}_{t-k},t-k,c)\in\mathbb{R}^{B\times N\times D}$. In our SRA, we set $m\leq n$, $k\geq 0$, and $0\leq(t-k)<t_{max}$or SiTs, $t_{max}=1$ and for DiTs, $t_{max}=1000$. In practice, we truncate ($t-k$) to 0 if it is less than 0.. Hence, we can conduct self-representation alignment using the teacher’s output ${\mathbf{y}}_{\ast}$ and the student’s output transform $j_{\psi}({\mathbf{y}})\in\mathbb{R}^{B\times N\times D}$, where $j_{\psi}({\mathbf{y}})$ is a projection of the student encoder output ${\mathbf{y}}$ that through a lightweight trainable MLPs head $j_{\psi}$. Notably, this projection head can be discarded optionally after training, which enables SRA to provide guidance without altering any architecture in diffusion transformers.
@@
where $[i]$ is a patch index, $\mathrm{dist}(\cdot,\cdot)$ is a pre-defined distance calculation function, and $\zeta_{s}$, $\psi$ is the parameters of student diffusion transformer and the projection head. This objective is similar to that mentioned in REPA [81], except that we’re aiming to align with the diffusion transformer itself.
@@
2.3 EMA Teacher Network
In SRA, we do not need an off-the-shelf teacher to give the prior guidance. Meanwhile, using the output of the same model as the target to conduct supervision would cause training instability [12, 28]. Thus, we follow some works in self-supervised learning (SSL) [85, 8, 30] that build the teacher from past iterations of the student network using an exponential moving average (EMA) on the student weights. In specific, the updated role of EMA is $\zeta_{t}=\alpha\zeta_{t}+(1-\alpha)\zeta_{s}$, where $\alpha\in[0,1)$ is the momentum coefficient. We study different values of $\alpha$ in Section 3.2 and show that $\alpha=0.9999$ unchanged works surprisingly well in our framework. Moreover, we do not use other operations like clustering constraints [6, 7], batch normalizations [28, 67], and centering [8, 85] in SSL because we find that the training progress is already stable enough without applying these tricks.
@@ 3 Experiment
In this section, we seek to validate the effect of SRA through extensive experiments. In a specific, we mainly focus on answering the following questions:
@@
  • Dose SRA work on different baselines across different model sizes? (Figure 4)
  • Can SRA show comparable or superior performance against other methods that leverage either the representation training paradigm or the representation foundation model? (Table 3.3)
@@ 3.1 Experimental Setup
Implementation details. Unless otherwise specified, the training details strictly follow the setup in DiT [62] and SiT [56], including AdamW [54] with a constant learning rate of 1e-4, no weight decay, batchsize of 256, using the Stable Diffusion VAE [68] to extract the latent, and etc. We use ImageNet [18], where each image is preprocessed to the resolution of 256$\times$256 (denoted as ImageNet 256$\times$256), and follow ADM [19] for other data preprocessing protocols. For model configurations, we use the B/2, L/2, and XL/2 architectures introduced in the DiT and SiT papers, which process inputs with a patch size of 2. Additional experimental details, hyperparameter settings, linear probing details, and computing resources, are provided in Appendix 3.
Evaluation. We report Fréchet inception distance (FID [31]), sFID [59], inception score (IS [70]), precision (Pre.) and recall (Rec.) [44]. To ensure a fair comparison with previous methods, we also use the ADM’s TensorFlow evaluation suite  [19]with 50K samples and the same reference statistics. A detailed breakdown of each evaluation metric is included in Appendix D.
Sampler. For DiT [62], we use the DDPM sampler and set the number of function evaluations (NFE) as 250 by default. For SiT [56], we use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the NEF as 250 by default. The settings also align with those used in DiT and SiT.
Methods for Comparison. To compare performance, we utilize recent advanced methods based on diffusion models. In particular, We focus on three distinct categories of approaches: (a) Pixel diffusion: ADM [19], VDM+++++ + [40], Simple diffusion [35], CDM [33], (b) Latent diffusion with U-Net: LDM [69], and (c) Latent diffusion with transformers: DiT [62], SiT [56], SD-DiT [86], MaskDiT [84], TREAD [42], REPA [81], and MAETok [9]. Here in the diffusion transformers family, we choose to compare with DiT/SiT and their modifications with representation components involved but do not compare with works aiming at designing advanced architecture like lightningDiT [80] and DDT [74]. Detailed descriptions of each method can be found in Appendix E.
@@ 3.2 Component-Wise Analysis
Below, we provide a detailed analysis of the impact of each component. To save time, we use SiT-B/2 and train with SRA for 400K iterations for evaluation. Results are shown in Table 3.2.
Table 1: Component-wise analysis on ImageNet 256$\times$256 without classifier-free guidance (CFG). $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $[0,k_{max})$ denotes time interval $k$ is chosen randomly from 0 to $k_{max}$. PH. denotes whether to use the projection head.
@@
SiT-B/2 Baseline [56] 33.02 43.71
$106\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 34.85 40.58
$84\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.00 47.78
$104\xrightarrow{}$ 10 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.65 46.69
[…]
$62\xrightarrow{}$ 6 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 32.14 46.36
$82\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.31 50.13
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$33\xrightarrow{}$ 3 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 37.08 41.54
$83\xrightarrow{}$ 8 0.0 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 31.07 47.32
$83\xrightarrow{}$ 8 0.1 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.55 49.01
$83\xrightarrow{}$ 8 0.2 smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 30.70 47.72
$83\xrightarrow{}$ 8 0$[0,0.1)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.38 49.32
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.3)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.15 50.01
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{1}$ Gen’s mom. $\checkmark$ 29.06 50.08
$83\xrightarrow{}$ 8 0$[0,0.2)$ $\ell_{2}$ Gen’s mom. $\checkmark$ 29.54 49.25
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Student copy $\checkmark$ 35.71 42.18
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ SSL’s mom. $\checkmark$ 33.17 44.96
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\times$ 34.23 41.07
$83\xrightarrow{}$ 8 0$[0,0.2)$ smooth-$\ell_{1}$ Gen’s mom. $\checkmark$ 29.10 50.20
@@
Block layers for alignment. We begin by analyzing the effect of using different blocks of student and teacher for alignment. We observe that using the teacher’s last but not least few layers (e.g., 8) to regulate the student’s first layers (e.g., 3) leads to optimal performance. We assume that the first few layers need more guidance so they can catch semantically meaningful representation for subsequent generation. Meanwhile, there is a strong correlation between the quality of the representations of the teacher’s layers and the performance of the corresponding aligned student (we give the quantitative results and analysis in the latter Section 3.4). Based on these results, we set alignment layers as $83\xrightarrow{}$ 8, $166\xrightarrow{}$ 16, and $208\xrightarrow{}$ 20 for B, L, and XL models respectively as defaultFor DiTs, we set alignment layers as $73\xrightarrow{}$ 7, $146\xrightarrow{}$ 14, and $168\xrightarrow{}$ 16 by default because DiT’s discriminative behavior across layers is slightly different from SiT’s (see Figure 2 and Figure 7).
@@
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with lower noise level than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor[0,200)\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
Objective for alignment. We next compare three simple regression training objectives for alignment, including smooth-$\ell_{1}$, $\ell_{1}$, and $\ell_{2}$. Empirically, we find that all three objectives can lead to good performance and are stable during training. We adopt smooth-$\ell_{1}$ by default in future experiments.
Teacher network for alignment. In other generative learning studies, the EMA model is often employed solely for evaluation. However, as we need it to provide guidance during training, we study different updating methods. Here, we investigate several different strategies to build the teacher. First, we find that using teacher copied from a student would impair the performance, which testifies to our argument in Section 1. Next, we consider using the strategy used in self-supervised learning works [8, 85, 28] that updates the momentum coefficient from 0.996 to 1 during training. However, in our framework, this does not work well. Finally, we use momentum coefficient of 0.9999 unchanged, which is commonly used in other generative learning works [19, 62, 56] and find it is also suitable for our framework. Thus, we set the momentum coefficient as 0.9999 in future experiments.

### v5

@@ Abstract
Recent studies have demonstrated that learning a meaningful internal representation can accelerate generative training. However, existing approaches necessitate to either introduce an off-the-shelf external representation task or rely on a large-scale, pre-trained external representation encoder to provide representation guidance during the training process. In this study, we posit that the unique discriminative process inherent to diffusion transformers enables them to offer such guidance without requiring external representation components. We propose Self-Representation Alignment (SRA), a simple yet effective method that obtains representation guidance using the internal representations of learned diffusion transformer. SRA aligns the latent representation of the diffusion transformer in the earlier layer conditioned on higher noise to that in the later layer conditioned on lower noise to progressively enhance the overall representation learning during only the training process. Experimental results indicate that applying SRA to DiTs and SiTs yields consistent performance improvements, and largely outperforms approaches relying on auxiliary representation task. Our approach achieves performance comparable to methods that are dependent on an external pre-trained representation encoder, which demonstrates the feasibility of acceleration with representation alignment in diffusion transformers themselves. Refer to caption Figure 1: Left: Methods like MaskDiT and SD-DiT use an external representation task to guide diffusion transformer. Middle: Methods like REPA leverage an external representation foundation model as guidance. Right (our approach): We do not use any external representation component but still obtain such guidance through proposed self-representation alignment technique.
@@ 1 Introduction
Diffusion transformers [dit, sit, pixartalpha] and vision transformers [vit, swin, pvt] have held the dominant positions in visual generation and representation because of their scalability during pre-training [sora, sd3, dinov2, evaclip] and generalization capacity for downstream tasks [sam, groundingdino, infiniteyou, pixwizard].
@@
Recently, many works [maskdit, sddit, repa, tread] have explored leveraging representation components of vision transformers for diffusion transformer’s training and have shown that learning a high-quality internal representation can not only speed up the generative training progress but also improve the generation quality. These works either utilize the an external discriminative loss in representation learning (e.g., MAE’s [mae], IBOT’s [ibot]) shown in Figure 1(left) or leverage a large-scale pre-trained representation foundation model (e.g., DINOv2 [dinov2], CLIP [clip]) shown in Figure 1(middle) to give representation guidance for the diffusion transformer during the original generative training. However, the former methods require an external representation task, and the latter method relies on a powerful external pretrained encoder, which limits its usage scenarios when there are no good external encoders. Thereout, an intriguing yet underexplored problem has come to light: Can we obtain such representation guidance without external representation components?
@@
Figure 2: We empirically investigate the trend of latent representations in the original SiT-XL/2. Left: Using PCA [pca] to visualize the latent features, we observe that the features lead a process from bad to good when increasing block layers and decreasing noise level. Right: A similar trend can also be seen in the linear probing results on ImageNet. Investigation of DiT is provided in Appendix A, which leads to the similar results as SiT.
@@
Our observations: Different from the representation model that takes a clean image as input and then outputs semantically-rich feature, diffusion model often takes a noise latent as input and obtains cleaner one step-by-step. In other words, the generative mechanism by which the diffusion model operates can be generally considered as a bad to good process. Inspired by this behavior, we hypothesise that the representations in it also follow such a trend. To testify this, we perform an empirical analysis with recent diffusion transformers [sit, dit]. As shown in Figure 2(left), we first find out that the latent features in the diffusion transformer are progressively refined, moving from bad to good, as block layers increase and noise level decreases. Next, akin to the results in previous studies [repa, ddae], we observe that the diffusion transformer already learns meaningful discriminative representations as shown in Figure 2(right). Meanwhile, although the accuracy drops off after reaching a peak at about layer 20 because the model needs to shift away to focus on generating images with high-frequency details, the quality of the representations basically transfers from bad to good by increasing block layers and decreasing noise level. These indicate that the diffusion transformer gets roughly from a bad-to-good discriminative process when only generative training is performed.
@@
Our approach: The observations above motivate us to align the weaker representations in the diffusion transformer to the better ones in the original generative training, thereby enhancing the representation learning of the model without involving any external representation component. Driven from this inspiration, we present Self-Representation Alignment (SRA). As shown in Figure 1(right), SRA does not need any representation component; in essence, it aligns the output latent representation in earlier layer conditioned on higher noise to that in later layer conditioned on lower noise to achieve self-representation alignment. Meanwhile, in order to boost the performance, we obtain the target features from another model that shares the same architecture with the trainable model but updates weight by weighted moving average (EMA)***In the following part, we use term ‘student’ to represent the trainable model and ‘teacher’ to represent the EMA model in SRA for simplicity.. Furthermore, the student’s output latent feature is first passed through the projection layers to conduct a slight nonlinear transformation for better representation extraction and then aligned with the target feature output by the teacher. In a nutshell, our SRA can offer a flexible way to integrate representation guidance without external component needs and architectural modification.
@@
Finally, we conduct comprehensive experiments to evaluate the effect of SRA. After a series of component-wise analyses, we show that SRA brings significant performance improvements to both DiTs [dit] and SiTs [sit]. Moreover, our ablation study highlights the crucial role of internal representations in the success of SRA, which supports our central hypothesis: diffusion transformers can achieve representation alignment without external components.
@@
  • We analyze the representations in diffusion transformers and assume that the unique discriminative process makes it possible to achieve representation alignment without external components.
@@ 2.1 Preliminary: Training Object of DiT and SiT
As our method is built upon the denoise-based model (DiT) and flow-based model (SiT), to facilitate a more seamless introduction of SRA in the subsequent part, we now present a brief overview of the training object of these two types of models. We omit the class-condition here for simplicity. We also leave more detailed mathematical descriptions of these types of models in Appendix B.
@@
[…]
@@
where the marginal distribution of this ODE at time $p_{t}({\mathbf{x}})$ matches that of the forward process. To learn the velocity field, the model is trained to minimize the following loss function:
[…]
@@ 2.2 Self-Representation Alignment
Previous studies have demonstrated that learning good internal representations can both speed up the diffusion transformer’s training convergence. In SRA, our insight is aligning the student’s latent feature in the earlier layer conditioned on higher noise with that in the later layer conditioned on the lower noise of the teacher to conduct self-representation alignment without requiring any external representation components. As depicted in Figure 3, the goal of our simple training framework is to let the diffusion transformer not only predict noise or velocity but also align with better visual representations from itself. This operation thereby provides a simple way to enhance the representation learning in the diffusion transformer during generative training without the need to design complex representation regulation or introduce external representation foundation models.
@@
Formally, let $f$ be the trainable student model and $f_{\ast}$ be the teacher model. Considering input noise latent, timestep, and condition to be ${\mathbf{x}}_{t}$, $t$, and $c$. Then, we can obtain the student encoder latent output ${\mathbf{y}}=f^{m}({\mathbf{x}}_{t},t,c)\in\mathbb{R}^{B\times N\times D}$, where $B,N,D>0$ are the batchsize, number of patches and the embedding dimension for $f$, and $m$ denote the output from the $m^{th}$ layer in $f$. Similarly, the output of the teacher can be expressed as ${\mathbf{y}}_{\ast}=f_{\ast}^{n}({\mathbf{x}}_{t-k},t-k,c)\in\mathbb{R}^{B\times N\times D}$. In our SRA, we set $m\leq n$, $k\geq 0$, and $0\leq(t-k)<t_{max}$For SiTs, $t_{max}=1$ and for DiTs, $t_{max}=1000$. In practice, we truncate (tkt-k) to 0 if it is less than 0.. Hence, we can conduct self-representation alignment using the teacher’s output $\mathbf{y}$$_{\ast}$ and the student’s output transform $j_{\psi}({\mathbf{y}})\in\mathbb{R}^{B\times N\times D}$, where $j_{\psi}({\mathbf{y}})$ is a projection of the student encoder output $\mathbf{y}$} that through a lightweight trainable MLPs head $j_{\psi}$. Notably, this projection head can be discarded optionally after training, which enables SRA to provide guidance without altering any architecture in diffusion transformers.
@@
where $[i]$ is a patch index, $\mathrm{dist}(\cdot,\cdot)$ is a pre-defined distance calculation function, and $\zeta_{s}$, $\psi$ is the parameters of student diffusion transformer and the projection head.
@@
2.3 Teacher Network
In SRA, we do not need an off-the-shelf teacher to give the guidance. Meanwhile, using the output of the same model as the target to conduct supervision would cause poor results (both found in previous work [simsiam, byol] and our experiment). Thus, we build the teacher from past iterations of the student network using an exponential moving average (EMA) on the student weights. In specific, the updated role of EMA is $\zeta_{t}=\alpha\zeta_{t}+(1-\alpha)\zeta_{s}$, where $\alpha\in[0,1)$ is the momentum coefficient. Note that this is widely used in self-supervised learning (SSL) [ibot, dino, moco]. But we find the default updating role in SSL does not perform well here, instead, $\alpha=0.9999$ unchanged works surprisingly well in our framework. Experiments and analyses of different $\alpha$ settings are presented in Appendix F. Moreover, we do not use other operations like clustering constraints [clustcon1, clustcon2], batch normalizations [byol, byol-wob], and centering [dino, ibot] in SSL because we find that the training progress is already stable enough without applying these tricks.
@@ 3 Experiment
In this section, we mainly focus on answering the following questions:
@@
  • Does SRA work on different baselines across different model sizes? (Figure 4, and Table 3.3)
  • Can SRA show comparable or superior performance against methods that leverage either the external representation task or the external representation encoder? (Table 3.3, Table 3.3, Figure 6)
@@ 3.1 Experimental Setup
Implementation details. Unless otherwise specified, the training details strictly follow the setup in DiT [dit] and SiT [sit], including AdamW [adamw] with a constant learning rate of 1e-4, no weight decay, batchsize of 256, using the Stable Diffusion VAE [sd-vae] to extract the latent, and etc. For model configurations, we use the B/2, L/2, and XL/2 architectures introduced in the DiT and SiT papers, which process inputs with a patch size of 2. Additional experimental details, hyperparameter settings, linear probing details, and computing resources, are provided in Appendix 5.
Evaluation. We report Fréchet inception distance (FID [fid]), sFID [sfid], inception score (IS [is]), precision (Pre.) and recall (Rec.) [pre-rec]. To ensure a fair comparison with previous methods, we also use the ADM’s TensorFlow evaluation suite  [adm]with 50K samples and the same reference statistics. A detailed breakdown of each evaluation metric is included in Appendix D.
Methods for Comparison.
We compare with recent advanced methods based on diffusion models: (a) Pixel diffusion: ADM, VDM++++, Simple diffusion, CDM, (b) Latent diffusion with U-Net: LDM, and (c) Latent diffusion with transformers: DiT, SiT, SD-DiT, MaskDiT, TREAD, REPA, and MAETok. We give detailed descriptions of each method in Appendix E. Note that in the diffusion transformers family, we choose to compare with DiT/SiT and their modifications with representation components involved but do not compare with works aiming at designing advanced architecture like lightningDiT [lightningdit] and DDT [ddt].
@@ 3.2 Component-Wise Analysis
Below, we provide a detailed analysis of the impact of each component. We use SiT-B/2 and train with SRA for 400K iterations for evaluation. Results are shown in Table 3.2. We also provide some hyperparameter setting principles for more easily transferring SRA to new models in Appendix G.
Table 1: Component-wise analysis on ImageNet $\times256$ without classifier-free guidance. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. $m\xrightarrow{}n$ denotes aligning features from $m^{th}$ layer of the student with that from $n^{th}$ layer of the teacher. $0,k_{\max})$ denotes time interval $k$ is chosen randomly from 0 to $k_{\max}$. PH. denotes whether to use the projection head.
@@
SiT-B/2 Baseline [sit] 33.02 43.71
$6\xrightarrow{}10 0$$[0,0.2)$ 0.2 $\checkmark$ 34.85 40.28
$4\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 30.00 47.78
$4\xrightarrow{}10 0$$[0,0.2)$ 0.2 $\checkmark$ 30.65 46.69
$4\xrightarrow{}12 0$$[0,0.2)$ 0.2 $\checkmark$ 33.19 43.30
$2\xrightarrow{}6$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 32.14 46.36
$2\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 29.31 50.13
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 29.10 50.20
$3\xrightarrow{}3$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 37.08 41.54
$3\xrightarrow{}8$ 0.0 0.2 $\checkmark$ 31.07 47.32
$3\xrightarrow{}8$ 0.1 0.2 $\checkmark$ 29.55 49.01
$3\xrightarrow{}8$ 0.2 0.2 $\checkmark$ 30.70 47.72
$3\xrightarrow{}8$ 0[0,0.1)[0,0.1) 0.2 $\checkmark$ 29.38 49.32
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 29.10 50.20
$3\xrightarrow{}8$ 0[0,0.3)[0,0.3) 0.2 $\checkmark$ 29.15 50.01
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.1 $\checkmark$ 30.65 48.31
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 29.10 50.20
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.3 $\checkmark$ 29.28 49.72
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.4 $\checkmark$ 29.75 49.30
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\times$ 34.23 41.07
$3\xrightarrow{}8$ 0[0,0.2)[0,0.2) 0.2 $\checkmark$ 29.10 50.20
@@
Block layers for alignment. We begin by analyzing the effect of using different blocks of student and teacher for alignment. We observe that using the teacher’s last but not least few layers (e.g., 8) to regulate the student’s first layers (e.g., 3) leads to optimal performance. We assume that the first few layers need more guidance so they can catch semantically meaningful representation for subsequent generation. Meanwhile, there is a strong correlation between the quality of the representations of the teacher’s layers and the performance of the corresponding aligned student (we give the quantitative results and analysis in the latter Section 3.4). Based on these results, we set alignment layers as $3\xrightarrow{}8$, $6\xrightarrow{}16$, and $8\xrightarrow{}20$ for B, L, and XL models respectively as defaultFor DiTs, we set alignment layers as $3\xrightarrow{}7$, $6\xrightarrow{}14$, and $8\xrightarrow{}16$ by default because DiT’s discriminative behavior across layers is slightly different from SiT’s (see Figure 2 and Figure 8).
@@
Time interval for alignment. We then study the time interval (meaning the same as $k$ in Section 2.2) used for alignment. Here, we study fixed and dynamic intervals. Notably, we find that using the teacher’s features input with a lower noise level than the student’s leads to performance enhancement, and an interval value of 0.1 or the mean of 0.1 is optimal. We hypothesize that this is because lower noise levels can offer better representation guidance, but an excessively large time interval can hinder the model’s learning process, causing it to focus only on optimizing alignment loss at the expense of neglecting the generative aspects. As dynamic interval shows slightly better performance, we apply time interval as $0\sim 0.2$ in our feature experiments§§§For DiTs, we set time interval as $\lfloor[0,200)\rfloor$ by default since DiTs adopt the linear variance schedule with $t$ where $t\in\{0,1,2,...,999\}\cap\mathbb{Z}$..
Regularization coefficient for alignment. We also examine the effect of the coefficient $\lambda$ of self-alignment loss. Note that the results are all better than the baseline, and there is only a little variance for different $\lambda$. We choose 0.2 as default since it achieves overall better performance.